5400 kg m/s t= 67 s F= ______N


1
Expert's answer
2021-04-07T10:08:57-0400
"F=\dfrac{\Delta p}{\Delta t}=\dfrac{5400\ \dfrac{kgm}{s}}{67\ s}=80.6\ N."